# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ 4.3. Additional Metadata Issues
We observed 596 models that had their human-readable identifiers changed. Instances such as 222gate/Blurdus-7b-v0.1 $\rightarrow$ gate369/Blurdus-7b-v0.1 involve a change to the name of the model owner. Other instances, such as aaditya/openbiollm-llama3-70b $\rightarrow$ aaditya/Llama3-OpenBioLLM-70B involve a change to the model name itself. In both cases, manual effort is required to track down and map the original references to the new names, which can make dependency management tasks increasingly difficult. Often the only way to determine the correct mapping is by relying on Hugging Face to redirect the page associated with a former name to the one associated with the most current name. This suggests that Hugging Face maintains some internal mapping, but it is unclear how long this mapping persists or what might happen in the event of a name collision.
Table 3. Top-10 Most Common Model Names in the Ecosystem
@@
Model names, without the additional owner information, are not guaranteed to be unique. Similar to how repository forks have the same name (but different owners) on GitHub, fine-tuned or forked models may also have the same name yet different owners on Hugging Face. For this reason, developers often specify models using the “owner/model” naming convention described previously. If this happens consistently, ambiguities can be avoided. However, we observed 16,477 cases where developers referred to base models only by the model name, excluding owner information. For example, a developer might use “roberta-base” to refer to “FacebookAI/roberta-base.” While it could be reasonable to assume that this shorthand refers to the popular model, this may not be the case: in our dataset, there were 39 other models also named “roberta-base,” all with different owners. Table 3 shows the top-10 most common model names in the Hugging Face ecosystem and their frequencies. If one of these models were referred to by name only, it would be nearly impossible to determine which model was being referenced. For example, more than 5,600 models share the model name “ppo-LunarLander-v2.” This further stresses the need for widespread adoption of unique model identifiers.
4.3.3. Missing or nonsensical references.
There were 34,159 instances where we were unable to map a declared dataset to a dataset found publicly on Hugging Face. These consisted of 4,755 unique declarations. It is impossible to know with certainty, but such declarations likely consist of datasets that have been removed or made private, are from external sources, are actually dataset descriptors, contain typos, or, in some cases, are Hugging Face usernames.
@@
In other cases, some of the references were nonsensical: in our analysis of the graph of all collected components in the ML supply chain (described further in Section 5), we detected 684 cycles in total, including 675 (98.7%) trivial cycles in which a model declares itself as a base model, despite the fact that a model should not be able to be its own ancestor. All of this further motivates the need to use unique and standardized identifiers when referring to models and datasets, as well as a need for validating reference information.
4.3.4. Models as datasets.
@@
4.3.5. Shortcomings of Hugging Face.
While base models are declared by some model owners, in many cases, the precise relationship between a derived model and its base model is left ambiguous, at least in the available metadata. Hugging Face does provide a separate field for architecture relationships, but there is no standardized way to specify situations involving fine-tuning, quantization, or using outputs for training. This ambiguity can have significant consequences since, particularly in the licensing context, the nature of the relationship is important in determining how licensing terms should be applied. For example, it is permissible to use a model licensed under the llama3 license for fine-tuning but impermissible to use its outputs to train a competing model (Meta, 2024). Additional fields specifying the nature of relationships between models would be useful, but their inclusion would also introduce additional overhead for model owners.
@@
5. RQ1: The structure of the ML supply chain
Table 4. Top-10 Models by Likes
Model Likes License
runwayml/stable-diffusion-v1-5 10,987 creativeml-openrail-m
CompVis/stable-diffusion-v1-4 6,358 creativeml-openrail-m
stabilityai/stable-diffusion-xl-base-1.0 5,502 openrail++
meta-llama/Meta-Llama-3-8B 5,200 llama3
bigscience/bloom 4,647 bigscience-bloom-rail-1.0
meta-llama/Llama-2-7b 3,980 llama2
mistralai/Mixtral-8x7B-Instruct-v0.1 3,960 apache-2.0
stabilityai/stable-diffusion-2-1 3,778 openrail++
WarriorMama777/OrangeMixs 3,714 creativeml-openrail-m
meta-llama/Llama-2-7b-chat-hf 3,670 llama2
Table 5. Top-10 Models by Downloads
Model Downloads License
MIT/ast-finetuned-audioset-10-10-0.4593 677,161,532 bsd-3-clause
sentence-transformers/all-MiniLM-L12-v2 83,805,061 apache-2.0
facebook/fasttext-language-identification 53,517,573 cc-by-nc-4.0
google-bert/bert-base-uncased 52,724,375 apache-2.0
sentence-transformers/all-MiniLM-L6-v2 41,281,684 apache-2.0
openai/clip-vit-large-patch14 40,114,370 -
distilbert/distilbert-base-uncased 24,195,717 apache-2.0
openai/whisper-small 23,317,804 -
jonatasgrosman/wav2vec2-large-xlsr-53-english 18,212,315 apache-2.0
openai/clip-vit-base-patch16 16,999,162 -
Despite the challenges outlined in RQ0, we were still able to construct an ecosystem-level supply chain graph mapping the relationships between models and datasets. While we acknowledge that our supply chain graph represents only a subset of the total components in the full Hugging Face ecosystem, as noted previously, we are confident that we can still rely upon it to make observations about recurring trends and patterns.
To analyze the structure of the ML supply chain, we began by addressing the issues identified in RQ0 using different strategies depending on the issue. Where possible, we employed web-scraping techniques to pull metadata for models that were inaccessible through the Hugging Face API (Section 4.1). Models where even this analysis was impossible were excluded from the analysis. We standardized model/dataset names, preferring the unique Hugging Face ID (Section 4.3.2). Where naming ambiguities were present (multiple models with the same name, but different owners), we assumed the most popular model (Section 4.3.2). References that could not be mapped to existing models/datasets on Hugging Face were also excluded from analysis (Section 4.3.3). Models that were listed as datasets were excluded from the dataset analysis (Section 4.3.4).
@@ 5.1. Most popular models and datasets
Table 6. Top-10 Datasets by Likes
Dataset Likes License
fka/awesome-chatgpt-prompts 5,011 cc0-1.0
HuggingFaceFW/fineweb 1,539 odc-by
Open-Orca/OpenOrca 1,269 mit
OpenAssistant/oasst1 1,232 apache-2.0
gsdf/EasyNegative 1,128 other
Anthropic/hh-rlhf 1,092 mit
togethercomputer/RedPajama-Data-1T 1,025 -
Nerfgun3/bad_prompt 921 creativeml-openrail-m
tiiuae/falcon-refinedweb 765 odc-by
allenai/dolma 752 odc-by
Table 7. Top-10 Datasets by Downloads
Dataset Downloads License
hails/mmlu_no_train 16,146,384 mit
lighteval/mmlu 5,882,652 mit
argilla/databricks-dolly-15k-curated-en 4,671,897 -
ceval/ceval-exam 2,111,340 -
lavita/medical-qa-shared-task-v1-toy 1,792,638 -
lukaemon/bbh 911,736 -
haonan-li/cmmlu 633,608 cc-by-nc-4.0
chansung/requested-arxiv-ids-3 587,249 -
cais/mmlu 555,115 mit
allenai/ai2_arc 543,678 cc-by-sa-4.0
We begin with an overview of the landscape by describing the most popular models and datasets. We measure popularity in terms of engagement via the metrics provided by Hugging Face, likes and recent downloads. These metrics also serve our goal of capturing the current state of the ecosystem by highlighting what models and datasets are currently being engaged with. Tables 5 and 5 show the ten most engaged with models in our dataset by these metrics. Organizations such as OpenAI and Meta occupy several top positions. The models span a variety of use cases, including image and text generation.
Similarly, Tables 7 and 7 display the most engaged with datasets. The most common Task field is “Question Answering” for the most downloaded datasets and “Text Generation” for the most liked datasets. However, a variety of Data Modalities and Tasks are represented across both groups. For example, Data Modalities included both text and images, while Tasks included other specifics such as “Token Classification.” In fact, the most liked dataset by far, fka/awesome-chatgpt-prompt, is a repository of curated ChatGPT prompts rather than a training dataset for models, which suggests that it may be more prudent to examine download count to determine the most popular datasets. This is further supported by its correlation with the number of client projects on GitHub (Pepe et al., 2024).
@@ 5.2. Most depended upon datasets and base models
Our ability to evaluate the ML supply chain relies on the dependency information provided by model owners. On Hugging Face, this information is found in the model metadata, where model owners declare base models and datasets. Although many models on Hugging Face were not created independently, we note that only 117,245 out of 760,460 models (15.4%) declare any base model and, of these, most (111,568) declare only one base model. Consequently, the average number of declared base models is 0, as are the first quartile, median, and third quartile, as shown in Figure 3.
Figure 3. Models’ Number of Declared Base Models
Table 8. Top-10 Declared Base Models
Base Model Times Used Likes Downloads License
distilbert/distilbert-base-uncased 5,240 463 24,195,717 apache-2.0
stabilityai/stable-diffusion-xl-base-1.0 4,776 5,502 3,558,713 openrail++
runwayml/stable-diffusion-v1-5 2,438 10,987 3,790,115 creativeml-openrail-m
openai-community/gpt2 2,405 2,076 6,888,015 mit
unsloth/llama-3-8b-bnb-4bit 2,363 141 486,478 llama2
FacebookAI/xlm-roberta-base 2,169 526 6,196,387 mit
mistralai/Mistral-7B-v0.1 2,099 3,290 706,043 apache-2.0
google-bert/bert-base-uncased 1,576 1,665 52,724,375 apache-2.0
google-bert/bert-base-cased 1,472 235 5,191,167 apache-2.0
openai/whisper-small 1,469 180 23,317,804 apache-2.0
@@
In addition to base models, we investigate the datasets used for model training. We observe that 75,516 (9.9%) models declare at least one dataset. Of these, 64,343 (85.2%) declare a single dataset, 5,066 (6.7%) declare two datasets, and 1,637 (2.2%) declare three datasets. The number of datasets also exhibits long-tail behavior, with one model declaring 287 distinct datasets. However, the vast majority of models declare no dataset, with 0 declared models as the average and first three quartiles, as shown in Figure 4. We observe that this low percentage of declared datasets is particularly notable since, by their nature, nearly all ML models on Hugging Face should have been trained on at least one dataset, even if that dataset is not publicly available.
Figure 4. Models’ Number of Declared Datasets
Table 9. Top-10 Declared Datasets
@@
Table 9 describes the top-ten declared datasets. We observe no overlap with the top-ten most liked or downloaded datasets. This may be a consequence of the low declaration rate for datasets, a function of a dataset’s age, or the fact that popular “datasets” uploaded to Hugging Face may not be intended for model training, as observed above.
License declarations vary among the datasets. Half of the top-ten most declared datasets are made available under a license declared as “other.” One example from Google appears under six distinct licenses, including the Apache-2.0 license, several Creative Commons licenses, and an “other” license. While these are all listed in the dataset’s metadata, the dataset card’s field for licensing information simply says, “More Information Needed,” leaving it unclear as to how these licenses jointly apply to the dataset. This suggests that even very popular components can contain ambiguities that make the license compliance process more difficult.

### v2

@@ 5.4. Naming problems
We observed 596 models that had their human-readable identifiers changed. Instances such as 222gate/Blurdus-7b-v0.1 $\rightarrow$ gate369/Blurdus-7b-v0.1 involve a change to the name of the model owner. Other instances, such as aaditya/openbiollm-llama3-70b $\rightarrow$ aaditya/Llama3-OpenBioLLM-70B, involve a change to the model name itself. In both cases, manual effort is required to track down and map the new names to the original references, which can make dependency management tasks increasingly difficult. Often the only way to determine the correct mapping is by relying on Hugging Face to redirect the page associated with a former name to the one associated with the most current name. This suggests that Hugging Face maintains some internal mapping, but it is unclear how long this mapping persists or what might happen in the event of a name collision.
Table 3. Top 10 Most Common Model Names in the Ecosystem
@@
Model names, without the additional owner information, are not guaranteed to be unique. Similar to how repository forks have the same name (but different owners) on GitHub, fine-tuned or forked models may also have the same name yet different owners on Hugging Face. For this reason, developers often specify models using the “owner/model” naming convention described previously. If this happens consistently, ambiguities can be avoided. However, we observed 16,477 cases where developers referred to base models only by the model name, excluding owner information. For example, a developer might use “roberta-base” to refer to “FacebookAI/roberta-base.” While it could be reasonable to assume that this shorthand refers to the popular model, this may not be the case: in our dataset, there were 39 other models also named “roberta-base,” all with different owners. Table 3 shows the top 10 most common model names in the Hugging Face ecosystem and their frequencies. If one of these models were referred to by name only, it would be nearly impossible to determine which model was being referenced. For example, more than 5,600 models share the model name “ppo-LunarLander-v2.” While many of these are likely toy models or “Hello World” equivalents, others, such as distilbert-base-uncased-finetuned-emotion, might suggest the need for widespread adoption of unique model identifiers.
5.5. Missing or nonsensical references
There were 34,159 instances where we were unable to map a declared dataset (e.g., a dataset that a model reported relying on for training) to a dataset found publicly on Hugging Face. These consisted of 4,755 unique declarations. It is impossible to know with certainty, but such declarations likely consist of datasets that have been removed or made private, are from external sources, are actually dataset descriptors, contain typos, or, in some cases, are Hugging Face usernames.
@@ 5.5. Missing or nonsensical references
In other cases, some of the references were nonsensical. In our analysis of the graph of all collected components in the ML supply chain (described further in Section 6), we detected 684 cycles in total, including 675 (98.7%) trivial cycles in which a model declares itself as a base model, despite the fact that a model should not be able to be its own ancestor. All of this further motivates the need to use unique and standardized identifiers when referring to models and datasets, as well as a need for validating reference information.
5.6. Models as datasets
@@
5.7. Shortcomings of Hugging Face
While base models are declared by some model owners, in many cases, the precise relationship between a derived model and its base model is left ambiguous, at least in the available metadata. Hugging Face does provide a separate field for architecture relationships, but there is no standardized way to specify situations involving fine-tuning, quantization, or using outputs for training. This ambiguity can have significant consequences since, particularly in the licensing context, the nature of the relationship is important in determining how licensing terms should be applied. For example, it is permissible to use a model licensed under the llama3 license for fine-tuning but impermissible to use its outputs to train a competing model (Meta, 2024). Additional fields specifying the nature of relationships between models would be useful, but their inclusion would also introduce additional administrative burden for model owners.
@@ 5.7. Shortcomings of Hugging Face
5.8. Summary
In summary, we identified numerous challenges while gathering data from Hugging Face, ranging from incomplete model information to the inability to access existing data. We highlight the prevalence of these issues to alert others seeking to mine Hugging Face and to identify specific areas for improvement, both from those documenting their models on Hugging Face and for any future efforts to improve Hugging Face’s system.
6. RQ1: Analysis of the ML supply chain
Table 4. Top 10 Models by Likes
Model Pipeline Likes License
runwayml/stable-diffusion-v1-5 text-to-image 10,987 creativeml-openrail-m
CompVis/stable-diffusion-v1-4 text-to-image 6,358 creativeml-openrail-m
stabilityai/stable-diffusion-xl-base-1.0 text-to-image 5,502 openrail++
meta-llama/Meta-Llama-3-8B text-generation 5,200 llama3
bigscience/bloom text-generation 4,647 bigscience-bloom-rail-1.0
meta-llama/Llama-2-7b text-generation 3,980 llama2
mistralai/Mixtral-8x7B-Instruct-v0.1 text-generation 3,960 apache-2.0
stabilityai/stable-diffusion-2-1 text-to-image 3,778 openrail++
WarriorMama777/OrangeMixs text-to-image 3,714 creativeml-openrail-m
meta-llama/Llama-2-7b-chat-hf text-generation 3,670 llama2
Table 5. Top 10 Models by Downloads
Model Pipeline Downloads License
MIT/ast-finetuned-audioset-10-10-0.4593 audio-classification 677,161,532 bsd-3-clause
sentence-transformers/all-MiniLM-L12-v2 sentence-similarity 83,805,061 apache-2.0
facebook/fasttext-language-identification text-classification 53,517,573 $c$-by-nc-4.0
google-bert/bert-base-uncased fill-mask 52,724,375 apache-2.0
sentence-transformers/all-MiniLM-L6-v2 sentence-similarity 41,281,684 apache-2.0
openai/clip-vit-large-patch14 zero-shot-image-classification 40,114,370 -
distilbert/distilbert-base-uncased fill-mask 24,195,717 apache-2.0
openai/whisper-small automatic-speech-recognition 23,317,804 -
jonatasgrosman/wav2vec2-large-xlsr-53-english automatic-speech-recognition 18,212,315 apache-2.0
openai/clip-vit-base-patch16 zero-shot-image-classification 16,999,162 -
Despite the challenges outlined in RQ0, we were still able to construct an ecosystem-level supply chain graph mapping the relationships between models and datasets. While we acknowledge that our supply chain graph represents only a subset of the total components in the full Hugging Face ecosystem, as noted previously, we believe that it is appropriate to rely on it to make observations about recurring trends and patterns.
To analyze the structure of the ML supply chain on Hugging Face, we began by addressing the issues identified in RQ0 using different strategies depending on the issue. Where possible, we employed web-scraping techniques to pull metadata for models that were inaccessible through the Hugging Face API (Section 5.1). Models where even this analysis was impossible were excluded from the analysis. We standardized model/dataset names, preferring the unique Hugging Face ID (Section 5.4). Where naming ambiguities were present (multiple models with the same name, but different owners), we assumed the most popular model (Section 5.4). References that could not be mapped to existing models/datasets on Hugging Face were also excluded from analysis (Section 5.5). Models that were listed as datasets were excluded from the dataset analysis (Section 5.6).
@@ 6.1. Most popular models and datasets
Table 6. Top 10 Datasets by Likes
Dataset Declared task Size Likes License
fka/awesome-chatgpt-prompts Question Answering ¡1K 5,011 cc0-1.0
HuggingFaceFW/fineweb Text Generation 10B-100B 1,539 odc-by
Open-Orca/OpenOrca Text Classification 1M-10M 1,269 mit
OpenAssistant/oasst1 - 10K-100K 1,232 apache-2.0
gsdf/EasyNegative - ¡1K 1,128 other
Anthropic/$h$-rlhf - 100K-1M 1,092 mit
togethercomputer/RedPajama-Data-1T Text Generation 1M-10M 1,025 -
Nerfgun3/bad_prompt - ¡1K 921 creativeml-openrail-m
tiiuae/falcon-refinedweb Text Generation 100M-1B 765 odc-by
allenai/dolma Text Generation ¿1T 752 odc-by
Table 7. Top 10 Datasets by Downloads
Dataset Declared task Size Downloads License
hails/mmlu_no_train Question Answering -* 16,146,384 mit
lighteval/mmlu Question Answering 1M-10M 5,882,652 mit
argilla/databricks-dolly-15k-curated-en - 10K-100K 4,671,897 -
ceval/ceval-exam Text Classification 10K-100K 2,111,340 -
lavita/medical-qa-shared-task-v1-toy - ¡1K 1,792,638 -
lukaemon/bbh - 1K-10K 911,736 -
haonan-li/cmmlu Multiple Choice/Question Answering 10K-100K 633,608 $c$-by-nc-4.0
chansung/requested-arxiv-ids-3 - ¡1K 587,249 -
cais/mmlu Question Answering 100K-1M 555,115 mit
allenai/ai2_arc Question Answering 1K-10K 543,678 $c$-by-sa-4.0
* No size is provided since the dataset requires Python code execution
We begin with an overview of the landscape by describing the most popular models and datasets. We measure popularity in terms of engagement via the metrics provided by Hugging Face: likes and recent downloads. These metrics also serve our goal of capturing the current state of the ecosystem by highlighting what models and datasets are currently being engaged with. Tables 5 and 5 show the ten most engaged with models in our dataset by these metrics. Organizations such as OpenAI and Meta occupy several top positions. The models span a variety of use cases, including image and text generation.
Similarly, Tables 7 and 7 display the most engaged with datasets. The most common Task field is “Question Answering” for the most downloaded datasets and “Text Generation” for the most liked datasets. However, a variety of Data Modalities and Tasks are represented across both groups. For example, Data Modalities included both text and images, while Tasks included other specifics such as “Token Classification.” In fact, the most liked dataset by far, fka/awesome-chatgpt-prompt, is a repository of curated ChatGPT prompts rather than a training dataset for models, which suggests that it may be more prudent to examine download count to determine the most popular datasets. This is further supported by the download count’s correlation with the number of client projects on GitHub (Pepe et al., 2024).
@@ 6.2. Most depended upon datasets and base models
Our ability to evaluate the ML supply chain relies on the dependency information provided by model owners. On Hugging Face, this information is found in the model metadata, where model owners declare base models and datasets. Although many models on Hugging Face were not created independently, we note that only 117,245 out of 760,460 models (15.4%) declare any base model and, of these, most (111,568 — 95.2%) declare only one base model. Overall, most models do not declare any base models: with respect to the number of base models declared, the average, as well as the minimum, first quartile, median, and third quartile, are all 0. The maximum is 53 unique base models.
Table 8. Top 10 Declared Base Models
Base Model Times Used Likes Downloads Pipeline License
distilbert/distilbert-base-uncased 5,240 463 24,195,717 fill-mask apache-2.0
stabilityai/stable-diffusion-xl-base-1.0 4,776 5,502 3,558,713 text-to-image openrail++
runwayml/stable-diffusion-v1-5 2,438 10,987 3,790,115 text-to-image creativeml-openrail-m
openai-community/gpt2 2,405 2,076 6,888,015 text-generation mit
unsloth/llama-3-8b-bnb-4bit 2,363 141 486,478 text-generation llama2
FacebookAI/xlm-roberta-base 2,169 526 6,196,387 fill-mask mit
mistralai/Mistral-7B-v0.1 2,099 3,290 706,043 text-generation apache-2.0
google-bert/bert-base-uncased 1,576 1,665 52,724,375 fill-mask apache-2.0
google-bert/bert-base-cased 1,472 235 5,191,167 fill-mask apache-2.0
openai/whisper-small 1,469 180 23,317,804 speech-recognition apache-2.0
@@
In addition to base models, we investigate the datasets used for model training. We observe that 75,516 (9.9%) models declare at least one dataset. Of these, 64,343 (85.2%) declare a single dataset, 5,066 (6.7%) declare two datasets, and 1,637 (2.2%) declare three datasets. The number of datasets also exhibits long-tail behavior, with one model (tasksource/deberta-base-long-nli) declaring 287 distinct datasets. However, the vast majority of models declare no dataset, with 0 declared models as the average, minimum, and first three quartiles, while the maximum is 287 datasets. We observe that this low percentage of declared datasets is particularly notable since, by their nature, nearly all ML models on Hugging Face should have been trained on at least one dataset, even if that dataset is not publicly available.
Table 9. Top 10 Declared Datasets
@@
Table 9 describes the top ten declared datasets. We observe no overlap with the ten most liked or downloaded datasets. This may be a consequence of the low declaration rate for datasets, a function of a dataset’s age, or the fact that popular “datasets” uploaded to Hugging Face may not be intended for model training, as observed above.
License declarations vary among the datasets. Half of the ten most declared datasets are made available under a license declared as “Other.” One example from Google appears under six distinct licenses, including the apache-2.0 license, several Creative Commons licenses, and an “Other” license. While these are all listed in the dataset’s metadata, the dataset card’s field for licensing information simply says, “More Information Needed,” leaving it unclear as to how these licenses jointly apply to the dataset. This suggests that even very popular components can contain ambiguities that make the license compliance process more difficult.
@@ 6.3. Structure of the supply chain
As explained in Section 4.1, our mined dataset consists of over 760,000 models. Our final model supply chain graph, which features the reported dependency relationships between models, features 761,826 nodes (models) and 126,995 directed edges. We remove any edges related to trivial cycles discussed in Section 4.3.3 where a model reports itself as its own base. This leaves us with 126,320 directed edges. The number of models in the graph exceeds the number in the dataset because some models were declared as dependencies but were not observed on Hugging Face at the time of mining. These models might have been removed, made private, exist on another platform, or had a name change that we were unable to trace.
The constructed supply chain graph can provide valuable information to model owners and researchers. For example, we can determine which models in the Hugging Face ecosystem are the most relied upon, and thus for which models complete, accurate documentation, compliance checks, and security audits may be more critical.
6.3.1. Analysis of Removed Cycles
As noted above, we removed from the analysis 675 trivial cycles. Of these, 168 featured nodes that were connected to other models in the supply chain graph. There were 154 instances (92%) where a model reportedly depended on itself and was also reported as a base for a another model. There were 14 instances (8%) in which a model reportedly depended on itself and at least one other model. In most cases, these cycles likely represent typos introduced by the model owner, but they could potentially also be artifacts indicating iterative fine-tuning or some other additional training.
6.3.2. Analysis of graph degree and density
There were 635,966 models (83.5%) isolated from the rest of the graph with no incoming or outgoing edges, representing the vast majority. In other words, they have a degree of zero. The next largest portion of nodes, a cohort of 111,840 (14.7%), have degree one. (“Degree” here refers to the total count of incoming and outgoing edges.) Table 10 shows the breakdown for statistics for up to degree five. There was one model, distilbert/distilbert-base-uncased, that had an astounding degree of 5,240. For this model, all of these represented outgoing edges, meaning that 5,240 models recorded depending on distilbert/distilbert-base-uncased. Taking all this together, the graph’s density is very low ($2.18x10^{-7}$), reinforcing that most models do not record any model dependency relationships. We rely on networkx’s density function to calculate and return the graph density, using Equation 1.
[…]
Table 10. Nodes of Degree N in Model Graph
Degree* Node Count Percentage
0 635,966 83.5%
1 111,840 14.7%
2 6,643 0.9%
3 2,550 0.3%
4 1,392 0.2%
5 764 0.1%
* Degree refers to the total count of incoming and outgoing edges
There are 13,985 model nodes that have at least one outgoing edge, meaning they are depended upon by at least one other model. Almost half of these (7,032) have at least two outgoing edges. 620 model nodes have an outgoing degree of twenty or more, and 150 have an outgoing degree of one hundred or more. There are 116,576 model nodes that have at least one incoming edge, meaning they record depending on at least one base model. Only 5,673 of these records are based on two or more base models. All of this demonstrates, unsurprisingly, that models are more likely to rely on other models than they are to be relied on themselves. We also note that this shows that the machine learning ecosystem is upheld by more than just a handful of well-known and readily recognizable models, such as the 150 models with one hundred or more direct dependents.
6.3.3. Connected components (subgraphs)
There are 643,653 connected components (i.e., disjoint subgraphs) within our supply chain graph. Nearly all of these, 635,966 (98.8%), are trivial subgraphs including just a single node. An additional 3,550 subgraphs are made up of only two nodes. Table 11 breaks down the most common order frequencies for the subgraphs. From this, it is clear that the connected components constructed based on available documentation tended to be small, but there were 107 subgraphs that contained one hundred or more nodes.
Table 11. Subgraph Order Frequencies
Order Count
1 635,966
2 3,550
3 1,349
4 803
5 519
6.3.4. An analysis of the largest subgraph
The largest subgraph featured 28,879 nodes and 35,633 edges, giving it a density of $8.54x10^{-5}$. The majority of nodes in this graph, 22,099 (76.5%), have a degree of one. This is followed by 3364 (11.6%) nodes with degree two and 1235 (4.3%) nodes with degree three. There were multiple hubs within the supply chain that served to connect the various nodes. For example, there were 81 nodes that had a degree greater than fifty. These represent popular models that have many models fine-tuned or based on them. See Table 8 for details on the top ten such models. Ultimately, we see that the sparse nature that is representative of the entire Hugging Face supply chain graph also applies to its largest connected component.
6.4. Model lineages